OpenStudy (anonymous):

Solve The System Of Linear Equations x + y = 4 2x + 3y = 0

OpenStudy (cwrw238):

hint multiply first equation by -2 and add - this eliminates the variable x

OpenStudy (whpalmer4):

Another approach, as one of the equations can be easily manipulated to give you one of the variables in terms of the other, would be to do so and substitute that expression in the other equation. \[x+y=4\]\[x=4-y\]Now rewrite the second equation using \((4-y)\) wherever you see \(x\).\[2(4-y) + 3y = 0\]Solve that for \(y\), then find \(x\) from \(x = 4-y\).

OpenStudy (whpalmer4):

@cwrw238's method is called elimination, and this one is called substitution.